OpenStudy (anonymous):

square root of 224 simplified

OpenStudy (anonymous):

Well the calculator says 14.966629.

OpenStudy (domebotnos):

\[4\sqrt14\]

OpenStudy (unklerhaukus):

224 = 4 * 56 = 4 * 4 * 14 = 4^2 * 14 √224 = √ (4^2 *14) = 4√14